OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0. quantity of negative 9 plus or minus 3 square root of 5 all over 2 quantity of 9 plus or minus 3 square root of 5 all over 2 quantity of negative 9 plus or minus square root of 117 all over 2 quantity of 9 plus or minus square root of 117 all over 2

OpenStudy (anonymous):

\[x^2+9x+9\] This can be factored down to\[(x+3)(x+3)\] Which can be expressed as\[(x+3)^2\] Oh my god what is wrong with your teacher? Are they really that evil? Seriously, that problem is factored so easily and they're making you do the quadratic formula for it?

OpenStudy (anonymous):

Okay, hold on. \[-b \pm \sqrt\frac{ b^2-4ac }{ 2a }\] \[-9 \pm \sqrt\frac{ 81-(4)(1)(9) }{ 2 }\] \[-9 \pm \sqrt\frac{ 81-36}{ 2 }\] \[-9 \pm \sqrt\frac{ 45 }{ 2 }\] \[-9 \pm \sqrt 22.5\] \[x=-9 + \sqrt22.5\] and \[x=-9-\sqrt22.5\] What is this sorcery? I've gone over the answer multiple times and it seems to be correct yet not correct because it's not one of the answers.

OpenStudy (anonymous):

Taking off from where you went wrong. only b^2-4ac is under square root, not the whole fraction. \[-b \pm \sqrt{b^{2}-4ac} \] is numerator, 2a is denominator \[(-9 \pm \sqrt{45})/2\] square root of 45 simplifies to \[3 \sqrt{5}\] for a final answer of \[(-9 \pm 3\sqrt{5})/2\]

OpenStudy (anonymous):

thank you so much!!!!